The income statement and balance sheet for GATA, Inc. are provided below. (Note that the reported figures are expressed in thousands of dollars).
GATA, INC.
Balance Sheets as of December 31, 2016 and 2017
(in 000s)
  2016   2017   2016   2017
Assets Liabilities and Owners’ Equity
  Current assets Current liabilities
     Cash $ 500 $ 1,000    Accounts payable $ 500 $ 700
     Accounts receivable 500 400    Notes payable 900 700
     Inventory 900 600   
        Total $ 1,900 $ 2,000       Total $ 1,400 $ 1,400
Long-term debt $ 400 $ 600
Owners’ equity
   Common stock $ 600 $ 600
  Fixed assets    Accumulated retained earnings 500 600
      Net plant and equipment $ 1,000 $ 1,200        Total Equity $ 1,100 $ 1,200
  Total assets $ 2,900 $ 3,200 Total liabilities and owners’ equity $ 2,900 $ 3,200
GATA, INC.
2017 Income Statement
(in 000s)
  Sales $ 1,700
  Cost of goods sold 1,030
  Depreciation 100
  EBIT $ 570
  Interest paid 70
  
  Taxable income $ 500
  Taxes 200
  
  Net income $ 300
  
  Dividends $ 200
  Retained earnings 100
Requirement #1:

Calculate each of the following ratios for GATA Inc. for 2017 only and complete the table below.  (Do not round intermediate calculations. Round your answers to 4 decimal places (e.g., 0.9173 or 2.1648). Input profit margin, return on assets, and return on equity as percentages (e.g., if the Profit Margin = 0.1576, input your answer as 15.76).

2017 Industry Average
     a. Current ratio times 1.2031
     b. Quick ratio times 0.7814
     c. Total asset turnover times 0.5875
     d. Inventory turnover times 2.2791
     e. Total debt ratio times 0.3654
f. Equity Multiplier 1.5757
g. Times interest earned ratio times 6.9117
h. Profit margin % 20.58%
     i. Return on assets % 12.09%
     j. Return on equity % 19.05%
Requirement #2:
Use the ratios you calculated in Requirement #1 and the industry averages given in the table above to determine if GATA, Inc. is "above average" or "below average" in each of the following areas of financial performance.
Financial Leverage (Click to select)above averagebelow average
Asset Use Efficiency (Click to select)above averagebelow average
Operating Efficiency (Click to select)above averagebelow average
Liquidity (Click to select)above averagebelow average
Shareholder Return (Click to select)above averagebelow average
Requirement #3:
Based on DuPont analysis, how would you explain the difference in the return on equity (ROE) for GATA, Inc. vs. the industry as a whole in 2017 (i.e., why is the ROE for GATA higher/lower than the ROE for the average firm in the industry)?
Answer #1

a]

current ratio = current assets / current liabilities = 2,000 / 1,400 = 1.43 +

b]

Quick Ratio = (Cash equivalents + accounts receivables) / Current Liabilities  

Quick Ratio = (1,000 + 400) / 1400 = 1.00

c]

Total Asset Turnover = Sales / Average Total Assets

Average total assets = (opening total assets + closing total assets) / = (2,900 + 3,200) / 2 = 3,050

Total Asset Turnover = 1,700 / 3,050 = 0.56

d]

Inventory Turnover = COGS / Average Inventory

Average Inventory = (opening Inventory + closing Inventory) / 2 = (900 + 600) / 2 = 750

Inventory Turnover = 1,030 / 750 = 13.73